A parallelogram and a triangle have equal areas. The base and height of the parallelogram are 12 and 9. If the base of the triangle is 36, find its height.
A. 12 B. 9 C. 6 D. 3 E. None of these

Answer by richwmiller(17219)

you need to know the formula for the area of a parallelogram and the formula for the area of a triangle.
b*h=area parallelogram and 1/2 b*h=area of triangle
9*12=1/2*36*h
9*12=18*h
9*2*6=9*2*h
6=h
C. 6